OpenStudy (anonymous):

Trinh is catering a wedding that 112 people will attend. She is making two cakes, a chocolate cake and a carrot cake.5/8 of the guests will be served chocolate cake and the rest will receive carrot cake. How many guests will get carrot cake?

OpenStudy (anonymous):

the top three who answer this will get a medal

OpenStudy (hoblos):

5/8 of the guests will be served chocolate cake so (1- 5/8 = 3/8) will receive carrot cake 3/8 (112) = 42

OpenStudy (anonymous):

your the best

OpenStudy (hoblos):

XD